Tohban Report 2016-04-27
Jump to navigation
Jump to search
Tohban Reports | |
---|---|
Start Date: | 20 Apr 2016 |
End Date: | 27 Apr 2016 |
Tohban: | Jim McTiernan |
Tohban email: | jimm@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Quiet:
How many GOES flares occurred?
Flares above B, C, M, X class were 18 1 0 0
RHESSI flare list contains 0 over the time range 19-Apr-16 26-Apr-16
No overlapping RHESSI flares to count, cool down is still in process.
Memory Management
Decimation was active/vigorous for the full time period. All detectors are off. The SSR was at 6.7% at the end of the overnight passes.
Spacecraft Status
Currently cooling down, with power settings based on the 95% cooldown curve: The cold plates, 1 and 2 are at 134.3 and 132.9 K, down from 144.1 and 142.6 K on 2016-04-20.
Data Gaps
GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30
N_GAPS 1
GAP START TIME GAP END TIME GAP (SEC)
2016-04-21T05:55:00.000 -- 2016-04-21T06:00:00.000 300.00000
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4
N_GAPS 1
GAP START TIME GAP END TIME GAP (SEC)
2016-04-21T06:40:00.000 -- 2016-04-21T06:50:00.000 600.00000
NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES)
Detector issues
Detectors are off.
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation weight 6, no front decimation |
Night time data (fronts) | +/- 4 minutes |
Night time data (rears) | +/- 4 minutes |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |